# Building Symbiotic Artificial Intelligence: Reviewing the AI Act for a Human-Centred, Principle-Based Framework

### v2

@@ 4 Discussion
This section discusses the analysis of the literature concerning the challenges, trends and limitations faced while conducting the SLR. The release of the AI Act has shifted the focus towards a more human-centric approach even in the literature belonging to the more technical side of Computer Science. Individuals are not considered as mere users, but as human beings in all their dimensions, which must be included in the process of creating ethical AI systems [79]. Nevertheless, little indication is provided regarding the design patterns that can be employed to implement human control and oversight mechanisms in AI systems.
@@ 4.1 Grounding the Framework in the Literature
To better ground the framework introduced in Section 3.2, this section compares it with existing similar frameworks available in the literature. More precisely, three sets of principles, are taken as reference for the comparison [15, 41, 66]. The comparison among them and the proposed human-centered SAI principles is shown in Figure 4.
@@
In 2020, the Berkman Klein Centre published a report to inform policymakers, researchers, and industry stakeholders about the complex ethical challenges posed by AI [15]. It promotes the development and deployment of AI technologies that respect humans’ fundamental rights, values, and societal norms. The report maps 36 documents illustrating frameworks, guidelines, and principles proposed by governments, international organizations, academic institutions, and industries. It delineates eight principles to guide the creation of AI-based systems: Transparency and Explainability, Fairness and Non-Discrimination, Safety and Security, Accountability, Promotion of Human Values, Privacy, Professional Responsibility, and Human Control of Technology [15].
@@
The principles differ from the ones identified in this report in structure, dependencies, and subdimensions. From the SLR, a new underlying hierarchical structure emerged; for example, we consider Safety, Security, and Privacy on the same level and part of the Protection principle, whereas the report views Privacy as a separate concept. As opposed to the Berkman Klein Centre’s Report, the principle of Transparency in our proposal stands at a higher level than Explainability and Interpretability, which indicates that an AI system exhibits transparent behavior if it presents models that provide appropriate and understandable explanations that allow humans to interpret their outputs correctly. Similar considerations hold for the principle of Fairness: from the content analysis we performed, a fair behavior of an AI model is achieved through a non-discriminatory approach, respecting human rights and spreading legitimate information. In this context, Promotion of Human Values was not identified as a separate or standalone principle since it is the leitmotif of the framework: each principle aims at ensuring that human values, rights, and ethics are preserved and reflected in the interaction with products that feature AI. As the law in the European Union becomes more involved in the creation of AI systems, designers, developers and operators are taking on more responsibility, which means that they must comply with regulations “by design”. Concerning Human Control of Technology, it plays a crucial role in our review, and it was redefined as part of a greater principle, called Automation Level, in which we indicate that AI systems exhibit a spectrum of automated behavior with respect to human intervention, ranging from guaranteeing oversight to allowing complete control [15].
@@
Integrating HCI in the design and development of AI is a relatively new area of research. Ben Shneiderman, one of the pioneers of HCAI, focuses on three main properties–Trustworthiness, Safety, and Reliability–as the key to reach other principles and satisfy the desiderata of AI systems. Shneiderman stresses the role of a proper balance between automation and control to empower humans [66]. With respect to this view of HCAI, the framework proposed in this manuscript splits and incorporates the three values in different principles. For example, we argue that a system is trustworthy if it follows multiple principles. Similarly, a safe system is a system that protects users, chooses the right automation level for the task at hand, and behaves fairly [29]. Finally, reliability of a system is a property that is commonly known and used in Software Engineering, and requires the intelligibility of the system’s operations, guaranteed by our framework’s principles of transparency and lawfulness while ensuring that the system behaves in accordance with the purposes of the interaction [67].
@@
Concurrently with the release of the AI Act in the EU, in 2024, Stanford University published the Artificial Intelligence Index Report, which reports the current trends, tracks, and advancements in the context of AI. This report presents ten key takeaways that highlight how AI cannot fully replace humans in terms of dimensions like reasoning, common sense, and empathy. Industries continue to excel in AI research, with the United States leading in developing top models, though these advancements come with rising costs not only in terms of resources but also in ethics and law. In addition, the scientific progress in the context of Generative AI has proven to be a boost in productivity, but there is still a notable lack when it comes to evaluating the responsibility of AI models, which is addressed in the report in the Responsible AI chapter. It provides its definition and dimensions while stressing that the ethical misuse of AI-based systems and their fast worldwide spread is contributing to a rise in the number of incidents [41].
@@
Although the report focuses on AI’s geopolitical and economic trends, some of its takeaways reflect the principles that emerged from our SLR. For instance, Automation Level is significantly present in the first and seventh takeaways being, respectively “AI beats humans on some tasks, but not all” and “The data is in: AI makes workers more productive and leads to higher quality work”; as humans possess characteristics that cannot be fully reproduced by AI, therefore their oversight and control becomes necessary. The ninth and tenth takeaways “The number of AI regulations [in the United States] sharply increases” and “People across the globe are more cognizant of AI’s potential impact—and more nervous” relate to the Lawfulness principle, as it highlights how relevant the legal component influences the design and development of AI systems. The fifth takeaway reflects the property of Robustness, one of the meta-codes of our research; it states that “robust and standardized evaluations for Large Language Model (LLM) responsibility are seriously lacking” [40].
@@
The four key dimensions that characterize Responsible AI are: Privacy and Data Governance, Transparency and Explainability, Security and Safety, and Fairness. Referring to our framework, data governance is an intrinsic aspect of Liability since it refers to the establishment of policies to guarantee the security, quality, and ethical use of data. Explainability is a dimension of the Transparency principle, and security and safety are part of the Protection principle.
@@ 4.3 Limitations
This happens when the research papers selected for the review do not represent all the studies conducted on the topic. In fact, the personal biases of the reviewers can influence the selection and interpretation of papers. This was mitigated by deeply analyzing the selected papers to check their compliance with the objectives of the SLR and by involving multiple reviewers who independently assessed them.

### v3

@@ 4 Discussion
This section discusses the analysis of the literature concerning the challenges, trends and limitations faced while conducting the SLR. The release of the AI Act has shifted the focus towards a more human-centric approach even in the literature belonging to the more technical side of Computer Science. Individuals are not considered as mere users, but as human beings in all their dimensions, which must be included in the process of creating ethical AI systems [Worsdorfer2023Mitigating]. This aspect must not undermine the performance and flexibility of such systems, providing humans with context-aware solutions that are able to support and not replace them [Carnevale2024Humancentred]. Nevertheless, little indication is provided regarding the design patterns that can be employed to implement human control and oversight mechanisms in AI systems.
@@ 4.1 Grounding the Framework in the Literature
To better ground the framework introduced in Section 3.2, this section compares it with existing similar frameworks available in the literature. More precisely, three sets of principles, are taken as reference for the comparison [Fjeld2020Principled, Maslej2024Artificial, Shneiderman2022HumanCentered]. The comparison among them and the proposed human-centered SAI principles is shown in Figure 4.
@@
In 2020, the Berkman Klein Centre published a report to inform policymakers, researchers, and industry stakeholders about the complex ethical challenges posed by AI [Fjeld2020Principled]. It promotes the development and deployment of AI technologies that respect humans’ fundamental rights, values, and societal norms. The report maps 36 documents illustrating frameworks, guidelines, and principles proposed by governments, international organizations, academic institutions, and industries. It delineates eight principles to guide the creation of AI-based systems: Transparency and Explainability, Fairness and Non-Discrimination, Safety and Security, Accountability, Promotion of Human Values, Privacy, Professional Responsibility, and Human Control of Technology [Fjeld2020Principled].
@@
The principles differ from the ones identified in this report in structure, dependencies, and subdimensions. From the SLR, a new underlying hierarchical structure emerged; for example, we consider Safety, Security, and Privacy on the same level and part of the Protection principle, whereas the report views Privacy as a separate concept. As opposed to the Berkman Klein Centre’s Report, the principle of Transparency in our proposal stands at a higher level than Explainability and Interpretability, which indicates that an AI system exhibits transparent behavior if it presents models that provide appropriate and understandable explanations that allow humans to interpret their outputs correctly. Similar considerations hold for the principle of Fairness: from the content analysis we performed, a fair behavior of an AI model is achieved through a non-discriminatory approach, respecting human rights and spreading legitimate information. In this context, Promotion of Human Values was not identified as a separate or standalone principle since it is the leitmotif of the framework: each principle aims at ensuring that human values, rights, and ethics are preserved and reflected in the interaction with products that feature AI. As the law in the European Union becomes more involved in the creation of AI systems, designers, developers and operators are taking on more responsibility, which means that they must comply with regulations “by design”. Concerning Human Control of Technology, it plays a crucial role in our review, and it was redefined as part of a greater principle, called Automation Level, in which we indicate that AI systems exhibit a spectrum of automated behavior with respect to human intervention, ranging from guaranteeing oversight to allowing complete control [Fjeld2020Principled].
@@
Integrating HCI in the design and development of AI is a relatively new area of research. Ben Shneiderman, one of the pioneers of HCAI, focuses on three main properties–Trustworthiness, Safety, and Reliability–as the key to reach other principles and satisfy the desiderata of AI systems. Shneiderman stresses the role of a proper balance between automation and control to empower humans [Shneiderman2022HumanCentered]. With respect to this view of HCAI, the framework proposed in this manuscript splits and incorporates the three values in different principles. For example, we argue that a system is trustworthy if it follows multiple principles. Similarly, a safe system is a system that protects users, chooses the right automation level for the task at hand, and behaves fairly [ISO20199241210]. Finally, reliability of a system is a property that is commonly known and used in Software Engineering, and requires the intelligibility of the system’s operations, guaranteed by our framework’s principles of transparency and lawfulness while ensuring that the system behaves in accordance with the purposes of the interaction [Siegel2024Media].
@@
Concurrently with the release of the AI Act in the EU, in 2024, Stanford University published the Artificial Intelligence Index Report, which reports the current trends, tracks, and advancements in the context of AI. This report presents ten key takeaways that highlight how AI cannot fully replace humans in terms of dimensions like reasoning, common sense, and empathy. Industries continue to excel in AI research, with the United States leading in developing top models, though these advancements come with rising costs not only in terms of resources but also in ethics and law. In addition, the scientific progress in the context of Generative AI has proven to be a boost in productivity, but there is still a notable lack when it comes to evaluating the responsibility of AI models, which is addressed in the report in the Responsible AI chapter. It provides its definition and dimensions while stressing that the ethical misuse of AI-based systems and their fast worldwide spread is contributing to a rise in the number of incidents [Maslej2024Artificial].
@@
Although the report focuses on AI’s geopolitical and economic trends, some of its takeaways reflect the principles that emerged from our SLR. For instance, Automation Level is significantly present in the first and seventh takeaways being, respectively “AI beats humans on some tasks, but not all” and “The data is in: AI makes workers more productive and leads to higher quality work”; as humans possess characteristics that cannot be fully reproduced by AI, therefore their oversight and control becomes necessary. The ninth and tenth takeaways “The number of AI regulations [in the United States] sharply increases” and “People across the globe are more cognizant of AI’s potential impact—and more nervous” relate to the Lawfulness principle, as it highlights how relevant the legal component influences the design and development of AI systems. The fifth takeaway reflects the property of Robustness, one of the meta-codes of our research; it states that “robust and standardized evaluations for Large Language Model (LLM) responsibility are seriously lacking” [Maslej2024AI].
@@
The four key dimensions that characterize Responsible AI are: Privacy and Data Governance, Transparency and Explainability, Security and Safety, and Fairness. Referring to our framework, data governance is an intrinsic aspect of Protection since it refers to the establishment of policies to guarantee the security, quality, and safe use of data, while Explainability is a dimension of the Transparency principle.
@@ 4.3 Limitations
This happens when the research papers selected for the review do not represent all the studies conducted on the topic. In fact, the personal biases of the reviewers can influence the selection and interpretation of papers. This was mitigated by deeply analysing the selected papers to check their compliance with the objectives of the SLR and by involving multiple reviewers who independently assessed them.
@@
5 Conclusions
The spreading of AI in many contexts of our daily lives leads to the necessity of understanding its applications and potential improvements along with its limitations. Due to the rapid advancement of AI, ethical and legal aspects have to be carefully addressed when creating these systems with the objective to considering humans in all of their dimensions, and not as mere users not undermining their capabilities and avoiding any form of discrimination [ISO20199241210, Stahl2023Unfair]. Both academia and government are identifying solutions by promoting a new perspective called HCAI which represents the need to adopt a human-centred approach when creating AI systems. HCAI is the foundation to create SAI systems that foster a human-AI symbiotic relationship empowering humans rather than replacing them [Desolda2024Human]. To safeguard humans during the interaction process and the environment surrounding them, the EU enacted the AI Act regulating their usage employing a risk-based approach [EuropeanParliament2024Regulation].
This article aims to present the principles, resulted from a SLR, that can guide practitioners to create SAI systems while being compliant with the AI Act considering humans as the core of the development process. Confirming what [Desolda2024Humancentered] argued, our SLR highlights that building SAI systems requires a multidisciplinary and human-centred approach, focusing on supporting humans rather than replacing them [Shneiderman2022HumanCentered]. This implies ensuring AI’s ethical and lawful behaviour as one of the main crucial aspects to integrate in the interaction between humans and AI, highlighting that both parties should improve over time, learning from each other [Desolda2024Humancentered]. Although there are technical challenges that designers and developers could encounter when creating SAI systems, ethical and philosophical issues must be inevitably addressed during the process [lisiAcceptabilitySymbioticArtificial2024].
This research has proposed a framework for building SAI systems that comply with the AI Act, composed of four principles— Transparency, Fairness, Automation Level, and Protection—and three properties—Trustworthiness, Robustness, and Sustainability. Both principles and properties foster symbiotic relationship, each to different extents, while guiding practitioners in creating legally-compliant and ethical systems. The results of this study lay the groundwork for the creation of a holistic framework, composed of guidelines and success criteria, that can provide them with more complete and concrete support through both design and evaluation methods, practices, and patterns.